OpenStudy (anonymous):

Can someone me solving a linear differential equations, cos(x) (dy/dx) +ysen(x) =2xcos^(3)x I have a problem with the integrating facytor u(x) =sec(x) but n get e^sec^(2)x

OpenStudy (anonymous):

which of these methods did you try separation of variable Exact Equations Bernoullis equation variation of parameters wronskian equation/determinant you will definitely get a correct answer with Wronkian provided you are great at intergration

OpenStudy (turingtest):

\[\mu(x)=\exp\left(\int\frac{\sin x}{\cos x}\right)\]

OpenStudy (turingtest):

\[\int\frac{\sin x}{\cos x}dx\]\[u=\cos x\implies du=-\sin dx\]\[-\int\frac1udu=-\ln(\cos x)=\ln(\frac1{\cos x})=\ln\sec x\]\[e^{\ln\sec x}=\sec x\]try that ;)

OpenStudy (anonymous):

he did.

OpenStudy (anonymous):

Great it was like that, turning test!

OpenStudy (turingtest):

welcome! I always have to recheck integral tanxdx myself, a little confusing